### v1

# Spin-glass states generated in van der Waals magnet CrGeTe by alkali-ion intercalation

Abstract
Tuning magnetic properties in layered van der Waals (vdW) materials has captured a significant attention due to the efficient control of ground-states by heterostructuring and external stimuli. Electron doping by electrostatic gating, interfacial charge transfer and intercalation is particularly effective in manipulating the exchange and spin-orbit properties, resulting in a control of Curie temperature () and magnetic anisotropy. Here, we discover an uncharted role of intercalation to generate magnetic frustration. As a model study, we intercalate Na atoms into the vdW gaps of pristine CrGeTe (CGT) where generated magnetic frustration leads to emerging spin-glass states coexisting with a ferromagnetic order. A series of dynamic magnetic susceptibility measurements/analysis confirms the formation of magnetic clusters representing slow dynamics with a distribution of relaxation times. The intercalation also modifies other macroscopic physical parameters including the significant enhancement of from 66K to 240 K and the switching of magnetic easy-hard axis direction. Our study identifies intercalation as a unique route to generate emerging frustrated spin states in simple vdW crystals.
1 Introduction
The discovery of long-range magnetic order in layered van der Waals (vdW) materials down to individual monolayers [1, 2] has opened up new avenues for improving the understanding of fundamental concepts in condensed matter physics [3, 4, 5, 6, 7, 8, 9, 10, 11, 12], together with potential for future applications [13, 14, 15, 16, 17]. For instance, a clean interface can be achieved in heterostructures made out of vdW materials and these devices may then be deployed for efficient switching of magnetisation state by spin-orbit torques [13, 14], as well as for magneto-optical applications [15, 16, 17].
Among the vdW magnetic materials, the families of chromium halides (i.e. CrI, CrBr and CrCl), lamellar ternary chalcogenides (i.e. CrGeTe and CrSiTe), metal thiophosphates (MPS, M = Fe, Mn and Ni) and iron-based tellurides (i.e. FeGeTe and FeGeTe), have been intensively and extensively studied [18, 19, 20, 21, 22]. One common aspect shared by most of these materials is the magnetic ordering transition temperature, Curie temperature () for ferromagnets, lying below room temperature [14], with a few exceptions including FeGeTe [14] and FeGaTe [23]. In order to fully realise the potential of 2D magnetic materials for practical applications, a consequential drive is taking place in either discovering new materials with above-room-temperature , or to elevate of existing systems by external means. For the latter, taking CrGeTe (CGT) as an example, its of 66 K [24, 25] is reported to be significantly enhanced by electrostatic gating [26, 27], strain [28, 29] and intercalation [7].
The modification of magnetic properties in vdW systems has been primarily focused on parameters such as and magnetic anisotropy constants that arise from the exchange and spin-orbit interaction respectively. Electron doping alters the type and strength of the exchange interaction [31, 32], e.g. via the change of the Fermi level position for itinerant ferromagnets [33], and the number of carriers that play a leading role in carrier-mediated indirect Ruderman–Kittel–Kasuya–Yosida (RKKY) and double-exchange mechanisms for less-conductive magnets [26]. Doped electrons populate in electronic bands with a specific orbital character, modifying the spin-orbit property impacting on magnetic anisotropies [34]. These material properties represent the homogeneous nature of systems. By contrast, some important yet unanswered questions in researching vdW magnetic systems are: what arise when a vdW magnetic system is introduced with inhomogeneous carrier-doping, and what are resultant parameter changes that we can experimentally access to? To address this with a model case, we employ Na-ion intercalation into the vdW gaps of pristine CrGeTe where Na ions are inhomogeneously intercalated within the gaps. We actively use this nature to intentionally create disorder and study its consequence to magnetism.
While inhomogeneity might be considered negative since it breaks the periodicity/coherency of crystallinity in real space, it can however generate new physics that cannot be achieved otherwise. In magnetism, spin-glass states [35] emerge because of the presence of multiple local energy minima that the system freezes into, generated by inhomogeneity and frustration/competition between the ferromagnetic and anti-ferromagnetic exchange interactions. Hopping between the energy minima originates experimentally measurable magnetic responses that are time-, temperature- and magnetic-field-dependent [10]. Characteristic of the phase transition, the spin-glass has an order parameter defined by Edwards and Anderson [37], together with the critical temperature of its freezing, . The research field of spin-glasses is vast and diverse, ranging from canonical spin-glass states often observed in non-magnetic materials with randomly distributed magnetic impurities [38] to more complex spin-glassy behaviours in frustrated ferromagnets, called reentrant spin-glasses [39]. Spin glass formation in highly correlated systems is also known to herald yet more exotic behaviour such as high-temperature superconductivity in the cuprates [40]. The realisation of glassy magnetic states in vdW materials therefore suggests their significant potential for supporting new magnetic and electronic states, formed by tuning material properties via substitution, intercalation and exfoliation or heterostructuring.
In this study, we intercalated pristine CGT single crystals with alkali (sodium, Na)-ions to create magnetically-disordered Na-CGT. As a result, we discovered a frustrated magnetic state in the vicinity of 35 K by dynamic magnetic susceptibility measurements suggesting the nature of this transition to be emergence of spin-glassiness behaviour. Further quantitative analysis points towards Na-CGT hosting a cluster spin-glass state at low temperatures, a feature not present in pristine CGT. This emergent frustrated state is accompanied by a clear increase of electric conductivity, enhanced Tc to K and the sign reversal of two-fold magnetic anisotropy term that removes its out-of-plane magnetic easy axis and creates the magnetic easy plane in Na-CGT. Our study illustrates that vdW magnetic systems are highly tunable, not only for the exchange and spin-orbit coupling properties, but also now for magnetic frustration by intercalation techniques.
2 Results Discussion
2.1 Characterisation of Na-CrGeTe
Figure 1 (a b) depicts the CGT and Na-CGT systems, respectively, and illustrates the presence of sodium-ions in the vdW gaps. Single CGT crystals were intercalated with sodium ions via the ammonia-based method [41] (details in Methods section). Structural changes in the intercalated CGT are characterised using X-ray diffraction (XRD) and Raman spectroscopy. XRD results in Fig. 1 (c d) show that the intercalated sample preserves the original crystalline structure of CGT. The c-axis parameter of pristine CGT is 6.8 Å, which corresponds to the distance between the centre of adjacent layers. A downshift is observed for all 00l-peaks after intercalation, indicating an increase in the interlayer spacing by 0.04 Å. This small increase in interlayer spacing upon intercalation is also observed in other intercalated compounds [42]. Raman spectroscopy is also performed for a large number of spatial positions on the sample surface, and the peak structure is maintained, consistent with the indication from XRD. However, quantitative analysis of the three most significant Raman modes (, , and ) shows that there are statistically significant changes to phonon energy and lifetime with intercalation. There are no major changes to the three main Raman modes that arise from intralayer vibrations of CGT, as expected, as intercalation should not change the bonding within the layers themselves. The small changes in peak energy and width (which is determined by lifetime) are consistent with carrier doping [9, 8] (see Supplementary Information Fig.S2-S3).
Electrical transport measurements were performed in order to see the effect of carrier doping by the alkali atoms to the system. Figure 1(f) shows the temperature dependence of resistance for both CGT and Na-CGT. Unlike the standard semiconducting behaviour of temperature dependence in pristine CGT [25, 45], Na-CGT displays a semiconductor-like response down to 80 K, before showing a metallic-like resistance behaviour in the low temperature regime. While complex, this clear change in temperature dependence of resistance further emphasises that electron doping by the intercalation process was successful.
Figure 2 (a) shows temperature-dependent magnetisation along the ab-plane orientation in both CGT and Na-CGT, and demonstrates the impact of Na intercalation on magnetisation in the static regime. In contrast to of CGT 66 K [24], Na-CGT shows its magnetisation values maintained to a higher temperature up to , implying this as its . Similar enhancement in CGT has been reported by organic molecule intercalation (to 208 K) [7] and by pressure (to 250 K) [28] where in all cases metallic transport accompanies the enhancement. Since scales with the exchange coupling strength, it is possible to interpret in our sample that doped electrons enhance the exchange coupling strength and at the same time these less-localised electrons can contribute to transport properties. In pristine CGT, electrons are fully localised and the super-exchange mechanism that arises from virtual electron excitations between two neighbouring Cr ions via a Te ion with the 90 Kanamori-Goodenough mechanism [46]. Local Hund’s rules within both Te and Cr ions stabilise a ferromagnetic order. Through this process, electrons have to be virtually excited into the conduction band via the bandgap, the exchange energy gain by this process is marginal, hence low in CGT. Doping real electrons to Cr ions changes their ionic states from Cr into Cr, activating the double-exchange mechanism [47] that bolsters the exchange interaction. It may also be relevant that since Cr is a Jahn-Teller ion, its substitution for Cr will lead to local distortions and increased local anisotropy that potentially enhances any spin glass behaviour that we discuss below.
CGT is known to have uniaxial magnetic anisotropy with the out-of-plane easy axis [24, 25]. Magnetisation measurements as a function of magnetic field (- curves) are performed for Na-CGT and the results at 10 K for the ab-plane and c-axis orientations are shown in Fig. 2 (b). It is evident that after intercalation, spins prefer to align within the ab-plane compared to the c-axis in CGT, implying the sign change of the uniaxial anisotropy term upon intercalation. Figure 2 (c) shows the - curves of Na-CGT at 200 K, showing a clear magnetic response for both orientations with the sustained uniaxial anisotropy. As a proof of the long-range magnetic order, such a response disappears at 295 K where only a linear - trend is observed and shown in Fig.S5(d) Supplementary Information.
Electrically conducting behaviour in Na-CGT allows us to probe the magnetic response via magnetoresistance (MR) which was impossible in pristine CGT. We performed angular dependent MR measurements along the sample orientation of in-plane (ab-plane) to the out-of-plane (c-axis) direction, as shown in Fig. 2(d), showing the clear magnetic anisotropy easy axis (plane) from the c-axis to easy-plane in Na-CGT since the saturation field becomes larger for more out-of-plane measurements. Figure 2(e) captures this change in magnetic anisotropy for different temperatures (calculation of the saturation field is highlighted in Fig.S11(b)) of Supplementary Information where for c-axis () increases with decreasing temperature, and so as the difference in to the ab-plane (). A sizable change in MR for Na-CGT across a wide temperature range is shown in Fig. 2(f), which disappears at the temperature where the long range order of also vanishes in Fig. 2(a). This measurement reassures consistency in the new intercalated Na-CGT samples and reproducibility of experimental data by means of separate probing techniques.
2.2 Emergent spin-glass nature
Another striking property which emerges via the intercalation process is spin-glass like behaviour in Na-CGT, a feature absent in pristine CGT. It is important to note that spin-glass states have not been experimentally reported in other heavily studied 2D magnetic systems [19, 20], such as CrI, CrBr, FeGeTe, as well as CGT with carrier doping by organic molecules [7] and electrostatic gating [26, 27], and pressure application [28, 29].
Spin-glass behaviour in Na-CGT is identified as the divergence of magnetisation in temperature between the zero-field cooled (ZFC) and field-cooled (FC) protocols, at small magnetic fields. As shown in Fig. 3 (a), a bifurcation between the two cases is observed below a temperature at which a steep increase in magnetisation takes places. The magnetisation of the ZFC curve shows a defined broad peak, whereas in the FC protocol, the magnetisation keeps on increasing as the sample is further cooled. The local maximum is observed and we define this as [48, 49]. This bifurcation in the magnetisation is minimised as the size of magnetic field in increased (see Supplementary Information). Such splitting and irreversibility in the magnetic memory of the system are recognised as characteristic features which can arise from various phenomena, including spin-liquid, super-paramagnetic, disordered anti-ferromagnetic and spin or cluster glass states [10, 38].
In order to gain further insight in this feature potentially representing a spin-glass like state in Na-CGT, ac magnetic susceptibility measurements were performed, where [12]. Figure 3 (b) shows the real (in-phase), , component of the susceptibility, as a function of temperature at different frequencies ranging from 50 Hz to 5 kHz. A distinctive cusp-like anomaly is quite evident in at low temperatures, which matches to the same temperature region of branching-off observed in Fig. 3 (a). In the inset of Fig. 3 (b), the main feature of is a small frequency shift in the peak-maximum towards higher temperatures as the frequency is increased. The shift in peak position for a frequency window of Hz-kHz indicates slow dynamics in the system, and this is contrary to ferromagnetic and conventional systems where a similar frequency dependence of is found only in the MHz regime [10]. Figure 3 (b) also highlights that for a small magnetic field disappears at high temperatures. The imaginary part (out of phase), , of the ac susceptibility shown in Supplementary Information (Fig.S8), is almost 30 times smaller at the peak maximum than . increases from zero to a finite value near the expected spin-glass like transition and also displays a clear anomaly with an inflection point, which is related to the dissipative nature of the system due to slow magnetic relaxation [11, 38].
The observed spin-glass property has formed within the ferromagnetically ordered matrix background. This particular type of spin freezing below a ferromagnetic transition is referred to as the reentrant spin glass state [39]. Here, the fluctuation of local exchange interaction is sufficient enough to produce the local energy minima that stabilise frozen spin-glass states against the ferromagnetic order [52, 53]. Phenomenologically, this condition is met when the averaged exchange coupling strength for the th and th site spins <> and its local fluctuation average <> are comparable to each other [39], where < · · · > denotes an average over the entire spin system. In Na-CGT, this situation is likely established due to inhomogeneous Na intercalation that induces an electron-doping distribution. Since the exchange coupling strength is a function of carrier density between magnetic ions, such doping inhomogeneity brings about the fluctuation and hence local competition of the exchange interaction, a key ingredient for the spin-glass states. While such atomic scale fluctuation is not directly measured in our study, the clear spin-glass signature observed in our study strongly indicates its presence.
For more quantitative analysis, a series of analytical procedures is carried out with phenomenological scaling and power-law models [10, 54] to further inquest and interpret the data with in as a function of frequency. The Mydosh parameter [10], , is measured as the respective variation of peak-maximum position per frequency, , defined as, . Here, corresponds to the difference in experimental temperature values of peak-maximum at high and low frequencies; is the difference in excitation frequencies for the given value. K sets a criterion for differentiating between superparamagetic transition () and spin-glass behaviour () [10, 55, 56, 57], and a similar frequency-dependent shift in has not been reported for anti-ferromagnetic or ferromagnetic phases at low -values [10]. For Na-CGT, is experimentally determined to be 0.01 which clearly suggests its spin-glass nature, sitting between the boundary of canonical and cluster spin glasses[58, 59].
In the next step, the standard theory of dynamic scaling near the phase transition temperature for given excitation frequencies is employed [10, 54]. This theory relies on a power law dependence of relaxation time (experimental attempt time), , as , where is the characteristic relaxation time related to the microscopic flipping time for dynamical fluctuating entities, such as a single moment or a cluster. corresponds to the freezing temperature in the limit of , and is the critical exponent. The power law equation is simplified into a linear form using a logarithmic base, and is adjusted in order to get the best linearity [54], where the goodness of fit is determined from the highest -coefficient value. Figure 3 (c) shows the result of this fitting process of the power law. is found to be 33.7 K, and then using this value, a linear least-squares fitting method is utilised to extrapolate s and . The obtained value of for Na-CGT lies in the range ( to s) typical for cluster spin-glass systems [59, 60] compared to canonical spin glasses [61, 62] ( to s) where the fluctuation of individuals moments happens at a faster rate. This indicates slower spin dynamics in Na-CGT at the transition temperature and it is most likely due to the presence of interacting spin-clusters in the system instead of individual spins. The value of zv for typical systems with spin-glass behaviour is reported between 4 and 12 [59, 62], and indeed this is also the case for Na-CGT, providing another experimental evidence for the cluster spin-glass nature in Na-CGT.
For magnetically interacting clusters, a modified form of the Arrhenius law, known as the Vogel-Fulcher(VF) law [11, 54] is used instead, to analyse ac susceptibility results. The Arrhenius law describes the relaxation of non-interacting spins [10, 11], and it is highlighted in Supplementary Information (Fig.S9) that even though the data seem to fit an Arrhenius law, the fit would imply unphysical values for the relaxation time and activation energy, suggesting that a description in terms of non-interacting spins is inadequate. The phenomenological VF law for the excitation frequency, , takes the following form, , where the characteristic frequency, , is obtained using the characteristic relaxation time from the scaling law, with and being the activation energy and the Boltzmann constant respectively. The constant, , is known as the VF temperature and attributed to inter-cluster interaction strength, and it is constrained having a value between 0 K to [63]. Figure 3 (d) shows a linear fit to the experimental data, where the parameters obtained from the least-squares fitting from the slope and intercept yield K and K, respectively. The agreement of the VF law to the experimental data and a finite value of insinuates interaction among the clusters, as the ratio between and provides information about the interacting strength between the clusters [11, 54]. In the case of Na-CGT, , which lies in the intermediate regime therefore highlighting a finite interaction between the dynamic entities.
Finally, the Casimir-du Pré equations are utilised [10, 12], in order to describe the dependence of both real and imaginary parts of the ac susceptibility on the excitation frequency at a given temperature. Figures 4 (a) and (b) show good agreement between the experimental data for both real/imaginary parts and the modified equations (Equations S3 S4) for different temperatures. The slow dynamics, arising in the vicinity of 35 K, are described by the generalised Debye model that takes into account the entities (i.e. spin-clusters for Na-CGT) having finite interactions leading to a system having a distribution of relaxation times due to the presence of cooperative effects [12]. is plotted against at a single temperature value and for different frequencies as shown in Fig. 4 (c), and we fit them by the following Cole-Cole equation:
where represents the susceptibility in the limit of zero frequency, known as isothermal susceptibility; is the adiabatic susceptibility for the high frequency limit; is a dimensionless parameter to represent its relaxation nature with 0 being the case for single relaxation time and 1 being for a broad distribution of relaxation times. The fact that the fit curves show an arc centered away from the origin indicates the presence of a distribution of relaxation times, compared to a semi-circle appearing for the system with a single relaxation time [12]. The presence of such distribution can be attributed to magnetic clusters with different sizes and anisotropies, each experiencing different relaxation times in the whole system. Further quantitative analysis can be available in Supplementary Information. This series of analysis confirms that at low temperatures (in the vicinity of 35 K), intercalation of CGT by Na-atoms leads to the co-existence of a cluster spin-glass transition with a background ferromagnetic state. The analysis point towards a progressive freezing of spin clusters and slower dynamics.
3 Conclusion
In summary, we show the unique feature of Na intercalation processes to significantly modify the transport/magnetic properties of vdW magnetic material CGT. We discovered that the intercalated Na-CGT hosts emerging magnetic states, i.e. the spin-glass states that are absent in pristine CGT. Such a spin-glass phase is a reentrant spin-glass system that coexists with the ferromagnetically ordered background. Dynamic susceptibility measurements analysed by multiple models all suggest the nature of cluster spin-glasses where spins are interacting to each other to behave collectively with the distribution of relaxation timescales. This study sheds light on the unique feature of intercalation to generate magnetically frustrated systems by inhomogeneous carrier doping. This new way of tuning vdW materials might be effective beyond the field of magnetism, where inhomogeneity will play an active role for determining material ground states.
4 Methods
Direct vapor transport method was used to synthesise the single-crystalline flakes of CrGeTe. Pre-mixing with a molar ratio of of high-purity elemental Cr ( in chips), Ge ( in crystals), and Te ( in beads) was carried out and then sealed in a quartz ampule at high vacuum ( Torr). The ampule was loaded in a single-zone furnace with increasing the temperature at a rate of 2 K/min reaching up to 1273 K. The furnace was set to slow cooling with the rate of 5 K/hour down to 673 K in order to obtain a low defect crystal growth. Then, nature cooling process was used with switching off the furnace. From the flux buildup, the single crystal flakes were separated and then stored in an inert atmosphere in a glovebox.
CGT was de-gassed in vacuum to a pressure less than 10 mbar at 100C. Na and CGT were placed in a custom-made air-tight glass tube at C and submerged in liquid ammonia, with a 1:1 molar ratio of Na:CrGeTe. The solution was maintained for a period of approximately 72 hours to allow for the intercalation to occur.
X-ray diffraction was performed using a Philips X’pert MPD diffractometer with Cu K radiation ( Å) in reflection geometry. For the measurement on air-sensitive Na-CGT, the sample was loaded into a berylium dome holder with O-ring type seal inside an Ar glovebox for measurement. Raman spectroscopy was performed using a Renishaw In-Via microscope equipped with a 488 nm laser and a 2400 I/mm grating, with a x20 magnification. Samples were measured in an air-free glass sample holder, loaded in a glovebox.
A Quantum Design Magnetic Properties Measurement System (MPMS-3) was used for dc variable-temperature susceptibility and isothermal magnetisation measurements. Measurements were taken during warming after cooling in zero field (ZFC-W) and in the measuring field (FC-W), where the measuring field was 200 Oe. Isothermal magnetisation measurements were performed in an external magnetic field between 70 kOe. A Quantum Design 9 T Physical Properties Measurement System (PPMS) using the ACMS-II option was used for ac susceptibility. Measurements were taken in a static base field with an oscillating ac driving field, where we used a combination of either 10 Oe and 10 Oe, respectively, or 50 Oe and 15 Oe, respectively. ac measurements were performed with the driving field oscillating at frequencies of 50, 100, 300, 500, 1000, 2500, and 5000 Hz.
For dc susceptibility measurements where external field is applied perpendicular to the layers of the crystal. The crystal was held tightly in place between two pieces of cling-film within a plastic cap which was suspended in the centre of a straw. For the inplane dc susceptibility measurements, the crystal was attached to the flat longitudinal edge of a semi-cylindrical quartz rod by Kapton. For ac susceptibility measurements, the magnetic field was not applied along a particular axis.
The electrical transport and magnetotransport measurements are performed by the four-point measurement technique [64, 65], as to characterise the electronic properties of the sample. A Keithley 2450 source meter is used to apply the current to the outer most connections and contact free resistance is measured by the two inner electrical connections. A copper sample box is used, inside which the sample is mounted. The sample stage is then placed on a probe copper arm inside of a closed cycle helium cryostat (Oxford Instruments), with a variable temperature range of 5 - 300 K. The probe arm of the cryostat sits between the poles of an electromagnet, which can produce a homogeneous magnetic field in the sample region.
Acknowledgements
S.K. and H.K. acknowledge support from the UK Engineering and Physical Sciences Research Council (EPSRC) via EP/T006749/1. E.S.Y.A. would like to thank the EPSRC for studentship funding from the Centre for Doctoral Training in Advanced Materials Characterisation (EP/S023259/1). L.A.V.N-C acknowledges a scholarship EP/R513180/1 to pursue doctoral research from EPSRC. A.S. thanks JSPS Postdoctoral fellowship for research in Japan (P21777). Magnetic measurements at the University of Cambridge were made on the EPSRC Advanced Characterization Suite (funded under EP/M0005/24/1). E.J.G.S. acknowledges computational resources through CIRRUS Tier-2 HPC Service (ec131 Cirrus Project) at EPCC (http://www.cirrus.ac.uk) funded by the University of Edinburgh and EPSRC (EP/P020267/1). EJGS acknowledges the Edinburgh-Rice Strategic Collaboration Awards and the EPSRC Open Fellowship (EP/T021578/1) for funding support. A.G.L. acknowledges financial support from the Spanish Ministry of Science, Innovation and Universities under grant No. PRX19/452.
Author contributions statement
S.K. and H.K. conceived the idea of the project. I.V. and G.E. synthesised and provided pristine CGT crystals. E.S.Y.A. and C.A.H. intercalated Na-ions in the CGT crystals. E.S.Y.A. performed Raman and XRD experiments. L.A.V.N-C. conducted dc-magnetometry and ac susceptibility experiments with supervision of S.E.D. S.K. performed the electrical transport and magnetotransport experiments. Data analysis and interpretations were carried out by S.K., E.S.Y.A., L.A.V.N-C., H.K. and all other co-authors. S.K., E.S.Y.A., L.A.V.N-C., A.S. and H.K. wrote the manuscript with input from the other co-authors.
Supplementary Information for “Spin-glass states generated in van der Waals magnet CrGeTe by alkali-ion intercalation"
S1 X-ray Diffraction
In order to distinguish background peaks from CGT 00l diffraction patterns, a graphite reference was also measured in a beryllium dome. The peaks shown in Figure S1 (a) corresponds to graphite peaks from literature, with the exception of the peak at 44 which is present in both diffraction patterns. Hence, the peak at 44 can be attributed to the sample environment, as opposed to Na-CGT.
To rule out presence of some common impurities in pristine and intercalated CGT, calculated XRD patterns with Cr-Te compounds were compared with the experimental data. As shown in Figure S1 (b), there is no overlapping of significant peaks between CGT and Cr-Te compounds.
S2 Raman spectroscopy analysis
The analysis of the Raman spectroscopy data was performed using the Python 3 S[5] package SciPy S[6]. Using the curve_fit function, a Lorentzian peak-shape was fitted to each of the modes studied, defined as:
where is the centre of the peak, is the full width at half maximum (FWHM), and is proportional to peak height.
| Centre (cm) | FWHM (cm) | |||||
| Material | ||||||
| CGT | 135.17(11) | 109.85(8) | 233.09(18) | 2.6(2) | 3.5(6) | 3.4(7) |
| Na-CGT | 136.03(16) | 109.8(6) | 234.2(5) | 3.8(4) | 4.4(11) | 7.1(13) |
Raman spectra from map scans of the two samples, with spatial resolution 10 m, were studied quantitatively in this way. A fit example is shown in Figure S2. The spectra for Na-intercalated CGT exhibited significant background throughout the measured range, which was not present in the spectra for the pristine sample. To account for this, a 12-order polynomial was fit to the regions of the data where there were no peaks. This background was subtracted before peak fitting occurred.
For each spectrum in the map scans, the fit was performed to obtain and FWHM for each of the , , and peaks. This was used to calculate an average and standard deviation, which is tabulated in Table S1. The full distributions are shown in Figure S3.
From this table it can be seen that there is a statistically significant increase in peak position for the and modes; within error, the mode does not change with intercalation. This consistent, statistically-significant change demonstrates that intercalation has taken place. However, it is a surprising result. With intercalation of an organic molecule in CGT, the Raman peaks decreased in energy due to the addition of electrons to the anti-bonding orbitals which weakened the bonds S[7]; this is also common to electron-doped layered materials, for instance gated S[8] or intercalated graphite S[9]. We speculate that this may occur to increased thermal conductivity of Na-intercalated CGT, leading to better dissipation of the heat from the laser.
We also observe an increase in peak-width with intercalation. This is similar to the result for intercalated graphite S[9]. It likely results from the reduced lifetime of the phonon due to electron-phonon coupling, though an alternative explanation could be spatial inhomogeneity within the sample. We note that the degree of the peak broadening depends on the different modes, implying its complex origins including spatial disorder/defects and electron-phonon coupling. The intercalation of an organic tetrabutyl ammonium (TBA) molecule into CGT S[7] led to the emergence of a new phonon mode while no new mode emerges in our Na intercalated CGT. We attribute this difference that TBA has a Raman-active mode associated with its intra-molecular vibrations, whereas Na is a single atom.
S3 DC-SQUID Measurements
As highlighted in the main text in Figure 2(a), the - curve in Na-CGT shows a different behaviour compared to CGT. In Figure S4, the magnetisation response as a function of temperature is compared along the two orientations of out of plane (c-axis) and in-plane (ab-plane) for the Na-CGT sample. It is evident from this measurement that the magnetic ordering temperature has been enhanced and the value of goes to zero in the vicinity of 240 K (inset of Figure S4).
In Fig. S5, the dependence of magnetisation as a function of external magnetic field (- curves) is shown for various different temperatures. Figure S5 (a) shows - curves for Na-CGT at 10 K for the ab-plane and c-axis orientations. Along the ab-plane, a saturation magnetic moment of 3.05 /Cr is obtained whereas along the c-axis it is found to be 2.87 /Cr. It is clear from the experimental - curves that it requires less external magnetic field for the magnetisation to saturate along the ab-plane orientation than c-axis. Hence, it is evident from these measurements that a change in the magnetic anisotropy takes place after intercalation of Na-ions into CGT, and the spins prefer to align in the in-plane orientation compared to the out of plane orientation.
Figure S5 (b c) show the - curves of Na-CGT at the elevated temperature of 200 K where a clear magnetic response is observable, together with the uniaxial magnetic anisotropy. The magnetic response disappears at 295 K as shown in Fig. S5(d) where a linear trend is observed. This confirms that the ferromagnetic ordering phase transition to a paramagnetic phase takes place in Na-CGT.
Finally, in Figure S6, - curves for CGT and Na-CGT for two different orientations are compared at 10 K. It is clearly visible that the magnetic state of CGT is modified post intercalation and a change in magnetic anisotropy takes place.
S4 Magnetic Susceptibility
The temperature dependence of the magnetisation in Na-CGT with small external magnetic field (20 mT) for the out of plane orientation is shown Figure S7 (a) and the bifurcation of magnetisation is also seen here with different field cooling protocols (similar to ab-plane orientation shown in the main text. Figure S7 (b) (c) show - curves for external field of 100 mT and this highlights that as the magnetic field is increased from 20 mT to higher field values, the branching between zero field-cooling (ZFC) and field-cooling (FC) protocols starts to disappear.
As mentioned in the main text, the ac magnetic susceptibility is expressed in complex form as , where the real part ,, is the in-phase component and the imaginary part, , is the out of phase component S[10]. Figure S8 shows the temperature dependence of for different excitation frequencies, and even though the amplitude is smaller compared to the real part shown in Fig. 3(b) in the main text, the anomaly feature is clearly present in , supporting the spin-glass picture. In typical spin-glass systems, relative low amplitude of is suggestive of distribution of relaxation mechanism instead of a single relaxation time S[11].
Figure S9 shows the ln vs 1/ plot of extracted from the peak-maximum in measurements. We used the following Arrhenius-type equation S[10, 11] to extract the characteristic activation energy as well as timescale given by:
where and is Boltzmann constant. As pointed out in the main text, even though the experimental data can be fit using Eq.(S3), the best fit parameters of = 8035 K and are both unphysical. The relaxation mechanisms in Na-CGT can thus not be described in terms of non-interacting spins, suggesting the interacting nature of Na-CGT’s magnetic clusters S[11].
The Casimir-du Pré equations analysis method is utilised as mentioned in the main text S[10, 12]. These equations describe the dependence of both real and imaginary parts of the ac susceptibility on the excitation frequency at a given temperature, and are used in the description of slow magnetic relaxation that takes the form of a Debye model. In the Debye model, a system relaxes for an entity (polarisation or magnetisation) and assumes a non-interacting nature between the components of that entity. However, as shown above for Na-CGT, this entity belongs to clusters with finite interaction strength which makes it difficult to describe the experimental data by a Debye model and leads to the system having a distribution of relaxation times due to the presence of cooperative effects. A modification of the Casimir-du Pré equations is carried out to take into account the broad distribution of relaxation times, referred to as the generalised Debye model S[12]. A detailed description of this model can be found elsewhere S[12]. This introduces a parameter, , which determines the spread of relaxation times, with two extremes of zero being the case for single relaxation time and the unity being for a broad distribution of relaxation times. The frequency dependence of ac susceptibility, (), in the generalised Debye model takes the following form S[12]:
where represents the isothermal susceptibility and is the adiabatic susceptibility S[12]. The Casimir-du Pré equations are modified by considering the following distribution function of :
As such, the real and imaginary parts of the complex ac susceptibility become:
where is the average relaxation time. We use these equations to fit the experimental data of the real and imaginary parts of ac susceptibility (e.g. those shown in Fig. 4). The temperature dependence of and is extracted by the fitting and shown in Figs. S10 (a)-(b) respectively. There is a clear increase in the average relaxation time in the vicinity of 35 K. This coincides with the earlier findings of bifurcation of magnetisation in different field-cooling protocols and the shift of peak-maximum in ac susceptibility measurements and it signifies slower dynamics as the freezing of the clusters takes place. The Cole-Cole analysis can also enables to determine by the equation:
extracted by the three different approaches, i.e. , and Cole-Cole analysis, all show qualitatively similar temperature-dependent trends with comparable values, as shown in Fig. S10(b). The values between 0.2 to 0.8 further support the picture of multiple channels of the spin-glass relaxation presented in the main text. Finally, we show the calculated by extracted from the Cole-Cole analysis in Fig. S10(c). This visualises the statistical distribution of relaxation times for magnetic clusters in Na-CGT.
S5 Magneto-transport
Here we show technical details and further analysis of magnetoresistance(MR) measurements measured in our Na-CGT. Figure S11(a) shows the MR measurements for Na-CGT at 10 K while the magnetic field is swept from positive to negative and then negative to positive directions. A peak-shaped resistance change is observed upon sweeping the external magnetic field around zero. The MR value plotted in Fig. 2(f) is deduced by the relative resistance difference between the high field and zero field values. We believe that this is anisotropic magnetoresistance where the resistance change depends on the angle between the current direction and magnetic moment orientation. On this assumption, the resistance represents the moment orientation and can be used for analysis of magnetic switching behaviour. In order to more reliably identify the completion of magnetisation switching, we use the derivative form of MR as shown in Fig. S11(b). This approach is also effective to remove any contribution from linear MR components.
The magnitude of the uniaxial magnetic anisotropy can be estimated by the difference in saturation fields for the out of plane (OOP) and in-plane (IP) orientations, as shown in Figure S11 (c). The finite difference representing the presence of magnetic anisotropies is another proof of the presence of long-range magnetic order. The temperature dependence is good agreement with the estimated from different measurement techniques such as magnetometry.